OpenStudy (anonymous):

Jake drops a book out of a window from a height of 10 meters. At what velocity does the book hit the ground?

OpenStudy (anonymous):

from the eq. of motion ,we know.v^2 - u^2 = 2as where v = final velocity u= initial velocity a= acc. s= distance a.t.q v^2 -o^2 =( 2 * 9.8* 1o) so, v = 14 m/s

OpenStudy (anonymous):

Exactly^ :)

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

Thank you guys(:

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Which statement best compares the momentum of a 5-kilogram fish swimming at a speed of 10 meters/second and a 2-kilogram fish swimming at a speed of 10 meters/second?

OpenStudy (anonymous):

How about that one^?:P

OpenStudy (anonymous):

This is a question which deals with conservation of momentum. If we can assume by the wording of the question that this is an inelastic collision (absorption with no recoil of any kind), then: MV = mv Where: M = mass of the larger fish V = velocity of the larger fish m = mass of the smaller fish v = velocity of the smaller fish So, when you plug the given values into the equation, you get (5 kg)(1 m/s) = (1 kg)(x m/s) Divide by 1 on both sides (toughie) and you're left with? :)

OpenStudy (anonymous):

5? But the answers dont come out like that they ask who has more momentum the 5kg fish or the 2kg swim? get me?

OpenStudy (anonymous):

or if both or niether?

OpenStudy (anonymous):

Okay then well the answer will be the 5kg fish :)

OpenStudy (anonymous):

Gracias :)